The rod with negligible mass rests on points A and B. The body with mass 22kg hangs on the C end of the rod. Find the forces acting on points A and B, if AC = 50cm and BC = 10cm.


1
Expert's answer
2021-03-23T11:12:11-0400

"\sum \tau^A=0\to(0.5-0.1)F_B-0.5mg=0\to0.4F_B=0.5\cdot22\cdot9.8\to"


"F_B=269.5\ (N)"


"\sum \tau^B=0\to-(0.5-0.1)F_A-0.1mg=0\to-0.4F_A=0.1\cdot22\cdot9.8\to"


"F_A=-53.9\ (N)"